The following data were taken from the balance sheet of Nilo Company at the end of two recent fiscal years:

Current Year Previous Year
Current assets:
  Cash $480,300 $386,400
  Marketable securities 556,200 434,700
  Accounts and notes receivable (net) 227,500 144,900
  Inventories 730,000 420,900
  Prepaid expenses 376,000 269,100
  Total current assets $2,370,000 $1,656,000
Current liabilities:
  Accounts and notes payable
  (short-term) $458,200 $483,000
  Accrued liabilities 331,800 207,000
  Total current liabilities $790,000 $690,000

a. Determine for each year (1) the working capital, (2) the current ratio, and (3) the quick ratio. Round ratios to one decimal place.

Current Year Previous Year
1. Working capital $ $
2. Current ratio
3. Quick ratio

b. The liquidity of Nilo has   from the preceding year to the current year. The working capital, current ratio, and quick ratio have all  . Most of these changes are the result of an   in current assets relative to current liabilities.
Answer #1

Part (a)

Working capital = Current assets - Current Liabilities

Current ratio = Current Assets/ Current Liabilities

Quick ratio = (Current assets - Inventory - Prepaid Expenses)/ Current Liabilities

Current Year Previous year
Working Capital (2,370,000 - 790,000) = $1,580,000 (1,656,000 - 690,000) = $966,000
Current Ratio (2,370,000/790,000) = 3 times (1,656,000/690,000) = 2.40 times
Quick ratio [(2,370,000 - 730,000 - 376,000)/790,000] = 1.60 times [(1,656,000 - 420,900 - 269,100)/690,000] = 1.40 times

Part (b)

The liquidity of Nilo has INCREASED from the preceding year to the current year. The working capital, current ratio, and quick ratio have all INCREASED . Most of these changes are the result of an INCREASE in current assets relative to current liabilities.